OpenStudy (amtran_bus):

How does sin* 1/cos give tan sec

OpenStudy (amtran_bus):

I know 1/cos is sec but where does tan come from?

OpenStudy (amtran_bus):

never mind its cos^2

Nnesha (nnesha):

\[\tan = \frac{ 1 }{ \color{reD}{\cot} }\] cot = cos /sin so \[\tan = \frac{ 1 }{ \color{reD}{\frac{ \cos }{ \sin }} }\] multiply 1 with the reciprocal of the bottom fraction \[\tan =\frac{ \sin }{ \cos }\] lolol

Nnesha (nnesha):

ohh okay I was like what type of question is this :P lol